# DIP: Unsupervised Dense In-Context Post-training of Visual Representations

### v1

@@ Abstract
We introduce DIP, a novel unsupervised post-training method designed to enhance dense image representations in large-scale pretrained vision encoders for in-context scene understanding. Unlike prior approaches that rely on complex self-distillation architectures, our method trains the vision encoder using pseudo-tasks that explicitly simulate downstream in-context scenarios, inspired by meta-learning principles. To enable post-training on unlabeled data, we propose an automatic mechanism for generating in-context tasks that combines a pretrained diffusion model and the vision encoder itself. DIP is simple, unsupervised, and computationally efficient, requiring less than 9 hours on a single A100 GPU. By learning dense representations through pseudo in-context tasks, it achieves strong performance across a wide variety of downstream real-world in-context scene understanding tasks. It outperforms both the initial vision encoder and prior methods, offering a practical and effective solution for improving dense representations. Code available here
@@ 1 Introduction
Our goal is to learn dense image representations from unlabeled data for effective in-context scene understanding. In-context learning, which enables models to adapt to tasks without updating parameters, has seen remarkable success in large language models (LLMs) [7]. By providing a few examples within the input prompt, LLMs generalize effectively to new tasks. Inspired by this, recent efforts aim to bring similar in-context learning capabilities to vision-only models [3, 5]. A notable approach by Balazevic et al. [3] reformulates dense prediction tasks as nearest-neighbor retrieval problems using patch feature similarities, demonstrating strong performance, especially with limited data. However, while recent self-supervised Vision Transformers (ViTs) [3, 56] show promise for in-context scene understanding, they still fall short of matching LLMs’ in-context learning success. To bridge this gap, it is crucial to learn dense features that establish strong semantic correspondences between patches in test and training images.
@@
Most existing methods, whether training from scratch or post-training, rely on self-distillation, where the teacher is an exponential moving average of the student weights. These frameworks use architectures and objectives designed to promote dense nearest-neighbor retrieval skills. For instance, state-of-the-art methods enforce consistency between objects across images and views [46] or between patch rankings from two random views of the same image [58]. However, self-distillation often suffers from poor stability [82] and hyperparameter sensitivity [2]. Moreover, these approaches introduce additional complexity, such as differentiable sorting [58], RoI alignment units [58, 83, 29], Sinkhorn-Knopp optimization [83, 64], or learnable patch pooling [3, 46]. While improving performance, these components make the methods harder to interpret and tune, and less transparent in terms of their underlying mechanisms.
@@
In this work, we adopt a post-training approach due to it’s computational efficiency and effectiveness. Within this context, we address two key questions: (1) Are there simpler and more effective alternatives to existing self-distillation frameworks with complex architectures and objectives?, and (2) Beyond the base model, can we leverage other vision foundation models to achieve our goal?
@@
We propose a simpler post-training approach with an unsupervised learning objective explicitly designed for in-context scene understanding. Inspired by meta-learning, which learns general knowledge from diverse tasks to address unseen but related tasks, we automatically construct multiple in-context tasks from unlabeled data and “meta-train” the model on them. This enables the model to acquire transferable dense representations for efficiently solving new in-context tasks during downstream stages. Unlike prior methods, our approach eliminates complex training objectives and architectures, avoiding reliance on self-distillation frameworks.
@@
To achieve unsupervised post-training, we generate in-context pseudo-tasks that mimic the tasks encountered during downstream use. Each pseudo-task consists of support examples with segmentation masks and a query image to be segmented based on the support examples. The support set includes a positive example (sharing objects with the query) and “distractor” examples (likely unrelated). Our pseudo-tasks are analogous to episodes from meta-learning literature [75] with the difference that we do not use any manual annotations during the post-training stage. With each episode, the model is essentially rehearsing for the downstream in-context segmentation task.
@@
We leverage pretrained VFMs to construct these pseudo-tasks. Specifically, apart from the base model (i.e., DINOv2R) that we post-train, we also use an auxiliary pretrained generative model, namely Stable Diffusion (SD) [27]. SD provides high-fidelity, class-agnostic image segments in an unsupervised manner, while the base model retrieves candidate positive examples for each query image and assigns pseudo-labels to the extracted segments. Although SD was trained on weakly annotated (internet-scraped) image-caption pairs, we consider our post-training method as unsupervised, as neither it nor the used pretrained models rely on dense segmentation labels.
@@
  • Compared to DINOv2R, DIP produces more semantically coherent neighbors, leading to more accurate retrieval-based scene understanding, particularly in low-shot settings (see LABEL:fig:segmentation_comparison). Additionally, DIP outperforms recent state-of-the-art post-training methods in most retrieval-based tasks while being significantly simpler.
@@ 2 Related Work
Self-supervised learning. Self-supervised approaches can learn rich representations from vast amounts of raw data via pretext tasks that provide supervision signals from the data itself. The model is subsequently fine-tuned on a downstream task of interest with few annotated examples. Self-supervision is commonly applied by hiding or transforming the input data and training the model to predict the missing parts or correct them. We distinguish several broad categories of self-supervised objectives for images: contrastive objectives for contrasting representations between visually similar and dissimilar views [54, 11, 30], clustering-based objectives [1, 8, 23], self-distillation objectives for predicting similar representations for different views of the same image [9, 26, 6, 24, 12], reconstruction objectives for predicting pixel colors [31, 79] or features for masked image patches produced by a teacher network [82, 2, 39, 25, 55]. Recent self-supervised approaches make use of pretrained VFMs for distillation in cooperation with the self-supervised objectives [4, 17, 59, 36]. These methods focus on learning global image representations that are tailored to classification tasks. However, without full-finetuning or task decoders, they are less suitable for scene-understanding tasks relying on dense representations, such as semantic segmentation. We build DIP on contrastive objectives [54] equipped with a memory bank [30].
@@
Dense self-supervision. To learn localized image representations that can be rapidly adapted to scene understanding tasks, specific self-supervised objectives are required to mimic the downstream task [78, 77, 33, 71, 46, 10, 3, 34, 68]. Contrasting features from dissimilar individual pixels or pseudo-segments from the same image can significantly improve object detection performance [78, 77, 33, 34, 53]. Another line of approaches propose different clustering strategies to produce object-aware supervision signals to improve semantic segmentation performance [71, 83, 10]. Others extend patch-level contrastive learning across similar and dissimilar images [3] enforcing nearest-neighbor consistency [46, 28, 58]. While we have similar objectives as  [58, 3] attain strong in-context scene understanding skills, we devise a different self-supervised objective inspired from few-shot learning that makes use of automatically computed object segment pseudo-labels.
@@
Recent works build upon pretrained VFMs [83, 28, 58, 69, 41] casting their self-supervised training as a post-training stage for preparing the VFM for dense downstream tasks. Similarly, DIP aims to rapidly endow pretrained VFMs with in-context reasoning skills.
@@
Unsupervised semantic segmentation. These works aim for producing object segments without supervision via self-supervised training. Earlier works in unsupervised segmentation focused mostly on different forms of clustering objectives, with segmentations refined through consistency across augmented views of the same image [40, 38, 37, 13, 76]. Another line of works explores pre-trained features from self-supervised VFMs [28, 66, 45, 47, 76] or from a diffusion UNet encoder [73, 15]. The recent DiffCut [15] leverages recursive Normalized Cuts in the final attention layer of a pretrained diffusion UNet encoder to produce object segments without supervision. Here we make use of DiffCut to extract pseudo-segmentation maps from the dataset images to be used for our in-context post-training. Although DiffCut segments can be noisy and imperfect, they can be good proxies for actual object segments for our post-training strategy.
@@
Few-shot learning. The aim of few-shot learning is to devise models that can rapidly acquire new skills only from a limited set of labeled examples. To that end, few-shot learning methods are trained on a distribution of few-shot tasks (also called episodes) sampled from the training dataset, that are similar to the few-shot tasks encountered at test time. There are numerous few-shot learning strategies, among which the most common ones are the methods that learn to match and map a query example to a class label by accessing a memory bank storing the few labeled examples for that task [75, 50, 65], metric-learning based approaches that learn distance metrics from few samples [70, 72], gradient descent-based approaches that learn how to adapt a model quickly to a given few-shot task via a limited number of gradient descent updates [20, 62], and generative methods that predict the weights of a classifier [60, 22, 21]. Inspired by Vinyals et al. [75] we devise a self-supervised pretext objective consisting of multiple in-context scene understanding tasks, each with a set of support samples stored in a memory and a query image to be segmented based on the support samples. In contrast to few-shot learning, we do not use any human annotation during the post-training stage and instead generate the segmentation labels automatically.
@@ 3 Method
Our goal is to fine-tune a pretrained vision encoder, typically a ViT [18],
to produce dense features suitable for in-context dense prediction tasks framed as nearest-neighbor retrieval. Given a set of support images (i.e., training images) with semantic segmentation annotations and a query image (i.e., test image), we aim to ensure that patch-wise features extracted from the query image are highly similar (e.g., high cosine similarity) to those from support patches depicting the same object category, while being dissimilar to features of different objects.
To achieve this, we propose an unsupervised post-training approach (Fig. 1) that fine-tunes the encoder on dense nearest-neighbor retrieval tasks automatically generated from unlabeled data. By training on pseudo-tasks simulating real-world scenarios, the encoder learns robust features transferable to real in-context tasks.
In the following sections, we first detail our dense in-context post-training approach Sec. 3.1 and then describe the automatic process for constructing dense in-context tasks Sec. 3.2, including the pseudo-labeling of class-agnostic segments and the selection of support examples.
@@ 3.1 Dense In-Context training
Label Pre-processing.  To align the label resolution with patch-wise features, we patchify the one-hot pseudo-labels $Y_{s_{i}}$. Specifically, we divide $Y_{s_{i}}$ into patches matching the encoder’s patch size and average the one-hot labels within each patch, yielding $Y^{\prime}_{s_{i}}\in\mathbb{R}^{L\times C}$, where $C$ is the number of pseudo-classes.
@@ 3.2 Automatic Dense In-Context Task Construction
For the second step, we assign pseudo-labels to each segment using the self-supervised DINOv2R feature encoder. We compute the mean DINOv2R feature for each DiffCut mask by pooling dense features within the mask region. To group visually similar segments, we apply K-means clustering on these pooled features [^1] For COCO [48], clustering is performed on the full dataset; for ImageNet [44], clustering is performed on a random subset of 200,000 training images for efficiency.. The resulting clusters serve as pseudo-classes for annotating the DiffCut masks.
To assign a pseudo-class to each DiffCut mask, we first assign each pixel-wise DINOv2R feature the cluster ID of its closest K-means centroid. Then, for each DiffCut mask, we perform majority voting on all pixels within the mask and assign the most frequent cluster ID. This ensures robust and consistent pseudo-labels, even with noisy features.
Selecting Examples for In-Context Task.  To generate in-context tasks, we pair each query image with one “positive” support image and $K$-1 “distractor” support images.
To find positive images sharing similar visual content with a query, we use global image representations from the DINOv2R encoder. For each image, we retrieve its five nearest neighbors from the dataset using DINOv2R global features. This results in five positive pairs per dataset image. We then refine this list by retaining only pairs where the query and positive-support images share a common pseudo-class occupying more than $5\%$ of each image’s area. This ensures semantically meaningful and relevant positive pairs. The final refined list contains all possible query-positive support pairs for training.
@@ 4.1 Experimental setup
$\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$6060606065656565707070707575757580808080Fraction of DatamIoU (%)(a) PascalVOC $\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$262626262828282830303030323232323434343436363636383838384040404042424242Fraction of Data(b) ADE20K
@@
Compared methods.  We compare our method with state-of-the-art unsupervised learning approaches, including DINO [9], Leopart [83], TimeT [64], iBOT [82], CrOC [71], CrIBO [46], DINOv2R [56, 16], and NeCo [58]. NeCo, a recent method (to appear at ICLR’25), is the most related to ours, as it also post-trains DINOv2R for dense representation learning in in-context scene understanding tasks. However, our post-training methodology differs significantly: we explicitly train on in-context pseudo-tasks relevant to our target applications, while NeCo adopts a self-distillation-based approach.
Training setup.  We apply our unsupervised post-training method to DINOv2 [56] with registers [16] (DINOv2R), using ViT-S/14 and ViT-B/14 models. For each model, in-context pseudo-tasks are automatically generated using DiffCut along with the corresponding base model (e.g., DINOv2R ViT-B/14 is used to generate pseudo-tasks for post-training DINOv2R ViT-B/14; see Sec. 3.2). Except otherwise stated, post-training is conducted on the COCO dataset for 5 epochs with 1,000 pseudo-classes. We also explore post-training CLIP and MAE models with ViT-B/16 (in this case using pseudo-labels generated with DiffCut and DINOv2R ViT-B/14), as well as post-training on ImageNet. Our post-training procedure is computationally efficient, requiring 17 hours and 52 minutes on a single V100 GPU with ViT-B, and 8 hours and 37 minutes on a single A100 GPU with ViT-S. Pseudo-label generation for COCO, which is done only once and offline, requires 17 hours on a V100 GPU. More implementation details in Appendix A.
@@ 4.2 Retrieval-based scene understanding
Our approach demonstrates consistent performance improvements over DINOv2R on all datasets, both in-domain (COCO) and out-of-domain (ADE20K, PascalVOC, Pascal-Context, and Cityscapes), highlighting the generalization capability of our learned representations. Compared to NeCo, our DIP consistently improves over DINOv2R, which is not the case for NeCo, achieving higher mIoU improvements in most cases and a higher average performance improvement (see “Avg. Delta” in Tab. 2).
@@
Monocular depth prediction.  In Tab. 3, we report results for in-context monocular depth prediction on NYUv2, measured by RMSE. Both DIP and NeCo improve over DINOv2R, with our method achieving the best RMSE score. This is notable, as our in-context pseudo-tasks are designed for semantic segmentation, yet the learned representations generalize effectively to the distinct depth prediction task.
@@
Post-training other base models.  In Tab. 4, we evaluate in-context segmentation on ADE20K and PascalVOC using CLIP [61] and MAE [31] as base models, both before and after post-training with DIP. Our method consistently improves performance across all tested models, demonstrating its versatility. Notably, the improvements for MAE are of $+33.4$ mIoU points on PascalVOC. These substantial gains transform MAE from a model poorly suited for in-context segmentation into one with significantly improved capabilities, highlighting the effectiveness of our approach for enhancing dense feature representations.
@@
Comparison with supervised baselines
To automatically generate pseudo-segmentations, our post-training approach leverages Stable Diffusion (SD) features [27], trained on weakly annotated (internet-scraped) image-caption pairs. In Tab. 5, we compare DIP features against both these SD features and those from supervised encoders (SAM [43] and RADIOv2.5 [32]). Crucially, SAM requires stronger supervision (manual segmentation masks) than SD, while RADIOv2.5 distills multiple model types (SAM, DINOv2, and CLIP-like features). Thus, both baselines leverage more supervision than our method. Additionally, RADIOv2.5 requires significantly more training compute. Despite these advantages, Tab. 5 demonstrates that DIP surpasses RADIOv2.5 in in-context semantic segmentation, confirming our method’s effectiveness. SAM underperforms in this setting, consistent with prior findings [42]. Importantly, DIP achieves significantly better performance than the SD features it utilizes for pseudo-segmentation.
Additionally, we provide linear segmentation results demonstrating consistent improvement over DINOv2R and competitive results compared to NeCo in Tab. 8 in supplementary.

### v2

@@ Abstract
We introduce DIP, a novel unsupervised post-training method designed to enhance dense representations in large-scale pretrained vision encoders for in-context scene understanding. Unlike prior approaches using complex self-distillation architectures, our method trains the vision encoder using pseudo-tasks that simulate downstream in-context scenarios, inspired by meta-learning principles. To enable post-training on unlabeled data, we propose an automatic mechanism for generating in-context tasks that combines a pretrained diffusion model and the vision encoder. DIP is simple, unsupervised, and computationally efficient, requiring under 9 hours on a single A100 GPU. By learning dense representations through pseudo in-context tasks, it achieves strong performance across a variety of downstream real-world in-context scene understanding tasks. It outperforms both the initial vision encoder and prior methods, offering a practical and effective solution for improving dense representations. See code here
@@ 1 Introduction
Our goal is to learn dense image representations from unlabeled data for effective in-context scene understanding. In-context learning, which enables models to adapt to tasks without updating parameters, has seen remarkable success in large language models (LLMs) [7]. By providing a few examples within the input prompt, LLMs generalize effectively to new tasks. Inspired by this, recent efforts aim to bring similar in-context learning capabilities to vision-only models [3, 5]. A notable approach by Balazevic et al. [3] reformulates dense prediction tasks as nearest-neighbor retrieval problems using patch feature similarities, demonstrating strong performance, especially with limited data. However, while recent self-supervised Vision Transformers (ViTs) [3, 56, 75] show promise for in-context scene understanding, they still fall short of matching LLMs’ in-context learning success. To bridge this gap, it is crucial to learn dense features that establish strong semantic correspondences between patches in test and training images.
@@
Most methods, whether trained from scratch or post-trained, rely on self-distillation, where the teacher is an exponential moving average of student weights [26]. These frameworks use architectures and objectives designed to promote dense nearest-neighbor retrieval skills. For instance, state-of-the-art methods enforce object consistency across images and views [46] or between patch rankings from two random views of the same image [58]. However, self-distillation often suffers from poor stability [83] and hyperparameter sensitivity [2]. Moreover, these approaches introduce additional complexity, such as differentiable sorting [58], RoI alignment units [58, 84, 29], Sinkhorn-Knopp optimization [84, 64], or learnable patch pooling [3, 46]. While improving performance, these components make the methods harder to interpret and tune, and less transparent in terms of their underlying mechanisms.
@@
In this work, we adopt a post-training approach due to its computational efficiency and effectiveness. Within this context, we address two key questions: (1) Are there simpler and more effective alternatives to existing self-distillation frameworks with complex architectures and objectives?, and (2) Beyond the base model, can we leverage other vision foundation models to achieve our goal?
@@
We propose a simpler post-training approach with an unsupervised learning objective explicitly designed for in-context scene understanding. Inspired by meta-learning, which learns general knowledge from diverse tasks to address unseen but related tasks, we automatically construct multiple in-context tasks from unlabeled data and train the model on them. This enables the model to acquire transferable dense representations for efficiently solving new in-context tasks during downstream stages. Unlike prior methods, our approach eliminates complex training objectives and architectures, and avoids self-distillation frameworks.
@@
To achieve unsupervised post-training, we generate in-context pseudo-tasks that mimic the tasks encountered during downstream use. Each pseudo-task consists of support examples with segmentation masks and a query image segmented using the support examples. The support set includes a positive example (sharing objects with the query) and “distractor” examples (likely unrelated). Our pseudo-tasks resemble meta-learning episodes [76] with the difference that we do not use any manual annotations.
@@
Instead, we leverage pretrained VFMs to construct these pseudo-tasks. Specifically, apart from the base model (DINOv2R) that we post-train, we use an auxiliary pretrained generative model, Stable Diffusion (SD) [27]. SD provides high-fidelity, class-agnostic image segments in an unsupervised manner, while the base model retrieves candidate positives for each query and assigns pseudo-labels to the segments. Although SD is pre-trained on image-caption pairs, we consider our post-training unsupervised as we do not use captions to extract image segments from SD, and neither SD nor the base model rely on dense segmentation labels.
@@
  • Compared to DINOv2R, DIP produces more semantically coherent neighbors, leading to more accurate retrieval-based scene understanding, particularly in low-shot settings (see LABEL:fig:segmentation_comparison). Additionally, DIP outperforms recent state-of-the-art post-training methods in most retrieval-based tasks while being simpler.
@@ 2 Related Work
Self-supervised learning. Self-supervised methods learn rich representations from vast amounts of raw data via pretext tasks providing supervision signals from the data, later enabling fine-tuning on downstream tasks with minimal annotations. The main self-supervised approaches for images include: contrastive objectives that distinguish similar and dissimilar views [54, 11, 30], clustering-based objectives [1, 8, 23], self-distillation objectives that match representations across augmentations [9, 26, 6, 24, 12], and reconstruction objectives that predict pixel colors [31, 80] or masked patch features from a teacher network [83, 2, 39, 25, 55, 75]. Recent approaches use pretrained VFMs for distillation in cooperation with self-supervised objectives [4, 17, 59, 36]. These methods focus on learning global image representations for classification tasks. However, without full-finetuning or task decoders, they are less suitable for dense scene-understanding tasks like segmentation. We build DIP on contrastive objectives [54] equipped with a memory bank [30].
@@
Dense self-supervision. Learning localized image representations that can be rapidly adapted to scene understanding requires specialized self-supervised objectives that mimic downstream tasks [79, 78, 33, 71, 46, 10, 3, 34, 68]. One approach contrasts features from dissimilar pixels or pseudo-segments within images, improving object detection [79, 78, 33, 34, 53]. Others employ clustering strategies to produce object-aware supervision signals to improve semantic segmentation [71, 84, 10], while some extend patch-level contrastive learning across images [3] with nearest-neighbor consistency [46, 28, 58]. Although we share similar objectives with [58, 3] for in-context scene understanding, we introduce a novel self-supervised objective inspired by few-shot learning that leverages automatically computed object segment pseudo-labels.
@@
Recent works build upon pretrained VFMs [84, 28, 58, 69, 41] casting their self-supervised training as a post-training stage for preparing the VFM for dense downstream tasks. Similarly, DIP aims to rapidly endow pretrained VFMs with in-context dense reasoning skills.
@@
Unsupervised semantic segmentation. These works aim to produce object segments via self-supervised training. Earlier works employed clustering objectives, with segmentations refined via consistency across augmented views [40, 38, 37, 13, 77]. Other works explore pretrained features from self-supervised VFMs [28, 66, 45, 47, 77] or diffusion UNet encoders [73, 15]. The recent DiffCut [15] leverages recursive Normalized Cuts in the final attention layer of a pretrained diffusion UNet encoder to produce object segments without supervision. We use DiffCut to extract pseudo-segmentation maps for our in-context post-training, as these segments - while sometimes noisy - serve as effective proxies for actual object segments in our framework.
@@
Few-shot learning. Few-shot learning develops models capable of rapid adaptation from limited labeled examples. These models are typically trained on few-shot tasks (or episodes) sampled from the training data, that simulate test-time conditions. Existing approaches include memory-based methods that match queries to labeled examples [76, 50, 65], metric-learning techniques that learn distance functions from few samples [70, 72], optimization-based methods that learn efficient adaptation through few gradient steps [20, 62], and generative approaches that predict classifier weights [60, 22, 21]. Inspired by Vinyals et al. [76], we devise a self-supervised pretext objective consisting of multiple in-context scene understanding tasks, each with a set of support samples stored in a memory and a query image to be segmented based on the support samples. In contrast to few-shot learning, we use no human annotation during post-training, generating labels automatically.
@@ 3 Method
Our goal is to fine-tune a pretrained vision encoder, typically a ViT [18], to produce dense features suitable for in-context dense prediction tasks framed as nearest-neighbor retrieval. Given a set of support images (i.e., training images) with semantic segmentation annotations and a query image (i.e., test image), we aim to ensure that patch-wise features extracted from the query image are highly similar (e.g., high cosine similarity) to support patches of the same object category, while being dissimilar to features of different objects.
To achieve this, we propose an unsupervised post-training approach (Fig. 1) that fine-tunes the encoder on dense nearest-neighbor retrieval tasks automatically generated from unlabeled data. Training on pseudo-tasks that simulate real scenarios, the encoder learns robust features transferable to real in-context tasks.
In the following sections, we first detail our dense in-context post-training approach (Sec. 3.1) and then describe the automatic process for constructing dense in-context tasks (Sec. 3.2), including the pseudo-labeling of class-agnostic segments and the selection of support examples.
@@ 3.1 Dense In-Context Training
Label pre-processing.  To align label resolution with patch features, we patchify the one-hot pseudo-labels $Y_{s_{i}}$. We divide $Y_{s_{i}}$ into patches matching the encoder’s patch size and average the one-hot labels within each patch, yielding $Y^{\prime}_{s_{i}}\in\mathbb{R}^{L\times C}$, where $C$ is the number of pseudo-classes.
@@ 3.2 Automatic Dense In-Context Task Construction
For the second step, we assign pseudo-labels to each segment using the self-supervised DINOv2R feature encoder. We compute the mean DINOv2R feature for each DiffCut mask by pooling dense features within the mask region. To group visually similar segments, we apply K-means clustering on these pooled features [^1] For COCO [48], clustering is performed on the full dataset; for ImageNet [44], clustering is performed on a random subset of 200,000 training images for efficiency.. The resulting clusters serve as pseudo-classes for annotating the DiffCut masks. To assign a pseudo-class to each DiffCut mask, we first assign each pixel-wise DINOv2R feature the cluster ID of its closest K-means centroid. Then, for each DiffCut mask, we perform majority voting on all pixels within the mask and assign the most frequent cluster ID. This ensures robust and consistent pseudo-labels, even with noisy features.
Selecting examples for in-context tasks.  To generate in-context tasks, we pair each query image with one “positive” support image and $K$-1 “distractor” support images.
[…]
@@ 4.1 Experimental Setup
$\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$606060656565707070757575808080Fraction of DatamIoU (%)(a) PascalVOC $\frac{1}{128}$$\frac{1}{64}$$\frac{1}{8}$$\frac{1}{1}$262626282828303030323232343434363636383838404040424242Fraction of Data(b) ADE20K
@@
Compared methods.  We compare our method with state-of-the-art unsupervised learning approaches, including DINO [9], Leopart [84], TimeT [64], iBOT [83], CrOC [71], CrIBO [46], DINOv2R [56, 16], and NeCo [58]. NeCo, a recent method (ICLR’25), is the most related to ours, as it also post-trains DINOv2R for dense representation learning in in-context scene understanding tasks. However, our post-training methodology differs significantly: we explicitly train on in-context pseudo-tasks relevant to our target applications, while NeCo adopts a self-distillation-based approach.
Training setup.  We apply our unsupervised post-training to DINOv2 [56] with registers [16] (DINOv2R), using ViT-S/14 and ViT-B/14 models. In-context pseudo-tasks are generated using DiffCut and the base model (e.g., DINOv2R ViT-B/14 is used to generate pseudo-tasks for post-training DINOv2R ViT-B/14; see Sec. 3.2). Except otherwise stated, post-training is on COCO for 5 epochs with 1,000 pseudo-classes. We also explore post-training CLIP and MAE models with ViT-B/16 (using pseudo-labels from DiffCut and DINOv2R ViT-B/14), as well as post-training on ImageNet. Our post-training is efficient, requiring 17h52min on V100 (ViT-B) and 8h37min on A100 (ViT-S). Pseudo-label generation for COCO, which is done only once and offline, requires 17h on a V100 GPU. More implementation details in Appendix A.
@@ 4.2 Retrieval-based Scene Understanding
Our approach shows consistent improvements over DINOv2R on all datasets, both in-domain (COCO) and out-of-domain (ADE20K, PascalVOC, Pascal-Context, and Cityscapes), highlighting the generalization capability of our representations. Unlike NeCo, our DIP consistently improves over DINOv2R, achieving higher mIoU and average performance improvements (see “Avg. Delta” in Tab. 2).
@@
Monocular depth prediction.  In Tab. 3, we report in-context depth results on NYUv2 (RMSE). Both DIP and NeCo improve over DINOv2R, with DIP achieving the best RMSE. This is notable, as our in-context pseudo-tasks are designed for semantic segmentation, yet the representations generalize well to depth prediction.
@@
Post-training other base models.  In Tab. 4, we evaluate in-context segmentation on ADE20K and PascalVOC using CLIP [61] and MAE [31] as base models, before and after DIP post-training. Our method consistently improves performance across all tested models, showing its versatility. MAE improves by +33.4+33.4+ 33.4 mIoU on PascalVOC. These gains transform MAE from weak to improved at in-context segmentation, highlighting our approach’s effectiveness in enhancing dense features.
@@
Comparison with supervised baselines  To automatically generate pseudo-segmentations, our post-training approach leverages Stable Diffusion (SD) features [27], trained on weakly annotated (internet-scraped) image-caption pairs. In Tab. 5, we compare DIP to SD and supervised encoder features (SAM [43] and RADIOv2.5 [32]). Crucially, SAM requires stronger supervision (manual segmentation masks) than SD, while RADIOv2.5 distills multiple model types (SAM, DINOv2, and CLIP-like features). Thus, both baselines leverage more supervision than our method. Additionally, RADIOv2.5 requires significantly more training compute. Despite these advantages, Tab. 5 shows DIP surpasses RADIOv2.5 in in-context semantic segmentation, confirming our method’s effectiveness. SAM underperforms in this setting, consistent with prior findings [42]. Importantly, DIP outperforms the SD features it uses for pseudo-segmentation.
Additionally, we report linear segmentation results with consistent gains over DINOv2R and competitive results compared to NeCo in Tab. 8 in supplementary.